Whirly Corporation’s contribution format income statement for the most recent month is shown below:

Total Per Unit
Sales (7,900 units) $ 237,000 $ 30.00
Variable expenses 158,000 20.00
Contribution margin 79,000 $ 10.00
Fixed expenses 55,200
Net operating income $ 23,800

Required:

(Consider each case independently):

1. What would be the revised net operating income per month if the sales volume increases by 100 units?

2. What would be the revised net operating income per month if the sales volume decreases by 100 units?

3. What would be the revised net operating income per month if the sales volume is 6,900 units?
Answer #1
Case 1 Case 2 Case 3
Units sold (7900+100)=8000 units (7900-100)=7800 6900
Sales (8000*30)=240,000 (7800*30)=234,000 (6900*30)=207,000
Variable expenses (8000*20)=160,000 (7800*20)=156,000 (6900*20)=138,000
Contribution margin 80,000 78,000 69,000
Fixed expenses 55200 55200 55200
Net operating income $24800 $22800 $13800.